# EvoChart: A Benchmark and a Self-Training Approach Towards Real-World Chart Understanding

### v1

@@ 5.3 Analysis
6 Conclusion
In this paper, we introduce EvoChart and EvoChart-QA: a novel approach for enhancing chart comprehension capabilities through self-training and iterative synthetic data generation, and a meticulously crafted real-world chart comprehension benchmark. We aim to provide a new avenue for real-world chart understanding through EvoChart and EvoChart-QA. Through extensive evaluation, we expose the limitations of existing VLMs in chart comprehension and validate the effectiveness of our EvoChart method across multiple datasets. In the future, we will continue to explore the integration of human-effort-free self-training methods with chart comprehension to further advance the field.
Figure 4: Overview of the EvoChart-QA Benchmark.
7 Appendix
7.1 Modified ChartQA
@@
7.2 EvoChart-QA Detailed Case
@@
7.3 EvoChart Action Space Types
@@
Is-Chart &
Is-Title-Clear
Drop
@@
Label-Value &
Value-Label
Drop
@@
Label-Visual &
Visual-Label
Drop
@@
7.4 EvoChart Method Question Template
@@
7.5 EvoChart-QA Source Websites

### v2

@@ 6 Conclusion
Acknowledgments
This work was supported by National Key Research and Development Program of China (2022YFC3303600), the Key Research and Development Project in Shaanxi Province No. 2022GXLH-01-03, National Natural Science Foundation of China (No. 62137002, 62293553, 62293554, 62450005, 62477036, 62293550, and 62306229), the Shaanxi Provincial Social Science Foundation Project (No. 2024P041), the Natural Science Basic Research Program of Shaanxi (No. 2023-JC-YB-593), the Youth Talent Support Program of Shaanxi Science and Technology Association (20240113), the China Postdoctoral Science Foundation (2024M752585).
Appendix
Modified ChartQA
@@ Acknowledgments
EvoChart-QA Detailed Case
@@
EvoChart Action Space Types
@@
Is-Chart & Is-Title-Clear Drop
@@
Label-Value & Value-Label Drop
@@
Label-Visual & Visual-Label Drop
@@
EvoChart Method Question Template
@@
EvoChart-QA Source Websites